OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (anonymous):

please help and show me how to do this step by step!

OpenStudy (anonymous):

M= y2 - y1/ x2 - x1 <<<gradient formula = 2-1 / 5-4 = 1 / 1 = 1 (4,1) y2-y1 = m(x2-x1)<<<find an equation formula y -1 = 1(x-4) y-1 = x - 4 y = x - 4 +1 y = x-3 <<<final equation

OpenStudy (anonymous):

idont get it:/ and i have 15 mor questinos of those

OpenStudy (anonymous):

ok, wait a second.

OpenStudy (anonymous):

questions*

OpenStudy (anonymous):

\[M = \frac{y2-y1}{x2-x1}\] That is the gradient/slope formula to find the gradient of the two points given. (4,1) and (5,2) (x1,y1) (x2,y2) so plug those in the formula, show me how you do it.

OpenStudy (anonymous):

2-1 5-2

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

no wait. ur 5-2 is wrong it's 5-4

OpenStudy (anonymous):

look carefully again.

OpenStudy (anonymous):

oh so its 0

OpenStudy (anonymous):

nono 5-4 = 1 not 0

OpenStudy (anonymous):

4-4 = 0 <<this would be equal to 0

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

look at the above steps that I did.